# Exploiting Spatial Diversity in Earth-to-Satellite Quantum-Classical Communications

### v1

@@ IV Coherent-State Transfer
Most of the existing classical optical communication systems rely on the transfer of modulated coherent states. To a large extent, the same can be said for quantum communication systems. For example, when a TMSV state is used for CV-QKD, the resulting EB CV-QKD protocol is equivalent to a Prepare-and-Measurement (PM) CV-QKD protocol, which relies on the transfer of small quantum-modulated coherent states [35]. Furthermore, the quantum-classical signaling adopted by the SQCC protocol proposed in [2] relies solely on the modulation of coherent states. As the first step towards the investigation of diversity-assisted SQCC over FSO fading channels we investigate the transfer of large (classically-modulated) coherent states (for the classical part of SQCC) and small (quantum modulated) coherent states (for the quantum part of SQCC) separately.
In general, our diversity-assisted coherent-state transfer aims to transfer the information encoded in the complex amplitude $\alpha_{\text{tar}}\in\mathbb{C}$ (the encoding amplitude) of a single-mode coherent state $|\alpha_{\text{tar}}\rangle$ (the encoded state) from Alice to Bob following the operational principles in Sec. II.2. More details on coherent states can be found in Sec. S-II of [38].
In this work, we investigate the usefulness of diversity in the transfer of both large coherent states (used for the classical part of communications) and small coherent states (used for the quantum part of communications). When large coherent states are used in classical communications, we assume that Alice uses the simplest Binary Phase-shift keying (BPSK) modulation by mapping each of her classical bits according to a classical alphabet $\alpha_{\text{tar}}\in\{-\alpha,+\alpha\}$ with $\alpha\in\mathbb{R}$ being a large number. Specifically, $\alpha_{\text{tar}}$ is drawn from a distribution whose PDF is given by
@@ IV.1 Quantum State Evolution
We use the Characteristic Function (CF) formalism to describe the quantum state evolution in the process of coherent-state transfer (more details can be found in [38]). In coherent-state transfer without using diversity, the coherent state $\smash{|\alpha_{\text{tx}}\rangle=|\alpha_{\text{tar}}\rangle}$ is sent through a lossy atmospheric channel with transmissivity $T$ and excess noise $\epsilon$. Using the beam-splitter description of the atmospheric channel (see Sec. B of [38]), we can express the CF of the output state $\rho^{(1)}$ that reaches the satellite receiver as
@@
When $M\geq 2$ subchannels are used in diversity-assisted coherent transfer, the transmitted state in each subchannel becomes $|\alpha_{\text{tx}}^{(M)}\rangle=|\alpha_{\text{tar}}/\sqrt{M}\rangle$ due to Alice’s equal splitting (recall Sec. II.2). After the transmission over subchannels and Bob’s diversity combining based on the configuration described in Appendix B, the CF of Bob’s final state can be readily expressed using the beam-splitter description detailed in Sec. S-III of [38].
For illustration purposes, in the following we explicitly present the CF of Bob’s final combined state in $M\times M$ diversity-assisted coherent-state transfer with $M$ ranging from 2 to 4. When $M=2$, the combining module consists of only one beam splitter (see Fig. A2(a) in Appendix B), and the CF of the final state $\rho_{\text{DC}}^{(2)}$ is given by
@@
When $M=3$, the combining module consists of two beam splitters (see Fig. A2(b) in Appendix B), and the CF of the final state $\rho_{\text{DC}}^{(3)}$ is given by
@@
When $M=4$, the combining module consists of three beam splitters (see Fig. A2(c) in Appendix B), and the CF of the final state $\rho_{\text{DC}}^{(4)}$ is given by
@@ IV.2 Performance Metric
where $\chi_{1}(\xi)$ ($\chi_{2}(\xi)$) denotes the CF of $\rho_{1}$ ($\rho_{2}$[48]. Naturally, the effectiveness of transferring any given encoded state $|\alpha_{\text{tar}}\rangle$ from Alice to Bob can be quantified by the fidelity between Bob’s final state and a scaled version of Alice’s encoded state $\smash{|\alpha_{\text{tar}}^{\prime}\rangle=\left|{\alpha_{\text{tar}}/{\langle\sqrt{T}\rangle}}\right\rangle}$. Specifically, the scaling factor $\smash{{1}/{\langle\sqrt{T}\rangle}}$ is applied to reflect the necessary operation in a real-world system – Bob has to scale his measurement results (using the available information of channel statistics) to compensate for the (fixed) mean channel loss. Under one realization of the $M$ fluctuating subchannel transmissivities, such a fidelity can be calculated as $\smash{\mathcal{F}\left(\rho_{\text{DC}}^{(M)},|\alpha_{\text{tar}}^{\prime}\rangle\langle\alpha_{\text{tar}}^{\prime}|\right)}$ using Eq. (21). The analytical expression of this fidelity is presented in Appendix C.
@@ V.1 Entanglement Distribution
In Fig. 2 we adopt all the settings in Appendix 3 and plot the scaled logarithmic negativity $E_{\text{LN}}^{(M)}$ and the RCI capacity $\mathcal{R}_{\text{C}}^{(M)}$ achieved by diversity-assisted entanglement distribution under different zenith angles ($\theta_{\text{z}}=0^{\circ}$, $\theta_{\text{z}}=30^{\circ}$, and $\theta_{\text{z}}=45^{\circ}$) in a real-world uplink quantum communication system. From the subfigures of Fig. 2, we first observe that the TMSV state with a larger quadrature variance $V_{\text{s}}$ (which equivalently indicates a stronger squeezing level) is more susceptible to fading – such an observation is consistent with the findings in previous works, (e.g., [49]). It can also be clearly seen that the incorporation of diversity improves the effectiveness of entanglement distribution, and the use of more subchannels leads to performance improvement in terms of both the logarithm negativity (Fig. 2(a)) and the RCI capacity (Fig. 2(b)). However, the performance improvement provided by adding subchannels is reduced as the number of subchannels increases. This “diminishing return” can be explained by Eq. (8) – indeed, when $M$ increases above a certain threshold, the non-Gaussian noise term $\frac{1}{M}\operatorname{Var}[\sqrt{T}](V_{\text{s}}-1)$ becomes negligible, making the total noise dominated by the term $(\langle\epsilon\rangle+1-T_{\text{eff}})$, which cannot be reduced by the use of diversity. Furthermore, the results in Fig. 2 also indicate that using diversity enables entanglement distribution over subchannels that are entanglement-breaking by themselves. For example, from Fig. 2(a) we can see that when $V_{\text{s}}=9$, it is impossible to distribute entanglement without using diversity under zenith angle $\smash{\theta_{\text{z}}=30^{\circ}}$ (see the $\smash{M=1}$ data point of the black dashed curve); however, the use of $2\times 2$ diversity enables the distribution of non-zero entanglement under the same channel condition (see the $\smash{M=2}$ data point of the black dashed curve).
(a) Scaled logarithmic negativity $E_{\text{LN}}^{(M)}$.
(b) RCI capacity $\mathcal{R}_{\text{C}}^{(M)}$.
Figure 2: (a) Scaled logarithmic negativity $E_{\text{LN}}^{(M)}$ and (b) RCI capacity $\mathcal{R}_{\text{C}}^{(M)}$ achieved by diversity-assisted entanglement distribution under different zenith angles $\theta_{\text{z}}=0^{\circ}$ (solid), $\theta_{\text{z}}=30^{\circ}$ (dashed), and $\theta_{\text{z}}=45^{\circ}$ (dotted) in a real-world uplink quantum communication system. All the subchannels are modelled via our phase-screen simulations. Different values of the quadrature variance $V_{\text{mod}}$ of Alice’s initial TMSV state are considered. Note that the $M=1$ data points represent the performance achieved without using diversity and are plotted as a benchmark.
@@ V.2 Coherent-State Transfer
We now examine the effectiveness of incorporating diversity for the purpose of coherent-state transfer within the context of SQCC. In addition to fading, the high mean loss over Earth-to-satellite channels can be very challenging for the faithful transfer of coherent states – indeed, our simulations indicate that the mean loss over a real-world Earth-to-satellite channel can be as high as ${\sim 40\,\text{dB}}$ (see Fig. A1 and Table A1 in Appendix 2 for details). In our investigation, we find that such a high mean loss essentially renders the faithful transfer of coherent states non-viable. It should be noted that this observation does not necessarily negate the possibility of classical communications, which focus on discriminating one coherent state from the other rather than the exact transfer of coherent states. Nevertheless, as in the case of coherent-state transfer, such a high loss will likely restrict the viability of the diversity-assisted discrimination of coherent states over fading channels (which has been investigated in [34]) as well. In order to demonstrate the usefulness of diversity as a means of fading mitigation in coherent-state transfer over other generic fading channels, we restrict ourselves to a regime with low mean subchannel losses and adopt a simple and flexible log-normal distribution (detailed in Appendix 4) to investigate the general behaviours of our diversity-assisted system over fading channels. Our focus on such a low-mean-loss regime makes our results (on average fidelity) more likely to be relevant to other FSO fading channels, such as those considered in [34] or satellite-based downlink channels that have a low mean loss and significant fading.
@@
Figure 3: Performance (average fidelity) of the diversity-assisted transfer of BPSK-modulated large coherent states with $\alpha$ ranging from $10$ ($100$ photons) to $50$ ($2500$ photons) achieved under 3 dB mean subchannel loss and 1 dB subchannel fading strength. All the subchannels are modelled using our log-normal model. Note that the $M=1$ data points represent the performance achieved without using diversity and are plotted as a benchmark.
@@
Figure 4: Performance (average fidelity) of the diversity-assisted transfer of small quantum-modulated coherent states with typical values of modulation variance $V_{\text{mod}}$ achieved under 3 dB mean subchannel loss and 1 dB subchannel fading strength. All the subchannels are modelled using our log-normal model. Note that the $M=1$ data points represent the performance achieved without using diversity and are plotted as a benchmark.
@@ VI Discussions
We have confirmed the usefulness of diversity for both entanglement distribution and coherent-state transfer over fading channels. Most straightforwardly, our findings indicate that employing diversity as a fading mitigation method can enhance the performance of all Gaussian quantum information protocols (including the standard CV-QKD) over fading channels (including Earth-to-satellite channels). As mentioned in Sec. II.1, in this work we have assumed that only the statistics of the transmissivity is known – in other words, the instantaneous channel transmissivity, sometimes referred to as the full Channel State Information (CSI) is not available to Bob. Interestingly, when the full CSI is available at Bob, we find that the diversity-assist system described in Sec. II.2 does not provide any non-negligible improvement in both entanglement distribution and coherent-state transfer under all our considered scenarios (results not presented).
@@ VII Conclusions
In this work, we explored the feasibility and effectiveness of utilizing spatial diversity as a means of fading mitigation in Earth-to-satellite CV quantum communications. We first determined a practical system model where spatial diversity can potentially be introduced to improve system performance within the context of satellite-based CV quantum communications. Based on our system model, we proposed a practical diversity-assisted scheme for the tasks of entanglement distribution and coherent-state transfer, and we analyzed the evolution of quantum states within these tasks. Using our numerical simulations for the realistic modelling of Earth-to-satellite channels, we perform a detailed performance investigation on diversity-assisted entanglement distribution, confirming the usefulness of diversity in Earth-to-satellite quantum communications. Within the context of SQCC over fading channels, our further results indicate that the transfer of both large (classically-modulated) coherent states (for the classical part of communications) and small (quantum-modulated) coherent states (for the quantum part of communications) can benefit from the use of diversity.

### v2

@@ IV Coherent-State Transfer
Most of the existing classical optical communication systems rely on the transfer of modulated coherent states. To a large extent, the same can be said for quantum communication systems. For example, when a TMSV state is used for CV-QKD, the resulting EB CV-QKD protocol is equivalent to a Prepare-and-Measurement (PM) CV-QKD protocol, which relies on the transfer of small quantum-modulated coherent states [36]. Furthermore, the quantum-classical signaling adopted by the SQCC protocol proposed in [2] relies solely on the modulation of coherent states. As the first step towards the investigation of diversity-assisted SQCC over FSO fading channels we investigate the transfer of large [^5] Note that, in this work we use the term “small coherent state” (“large coherent state”) to refer to a coherent state with photon number $\leq 10$ ($\geq 100$). We use the term ”coherent state” on its own to refer to a coherent state with any photon number. (classically-modulated) coherent states (for the classical part of SQCC) and small (quantum modulated) coherent states (for the quantum part of SQCC) separately.
In general, our diversity-assisted coherent-state transfer aims to transfer the information encoded in the complex amplitude $\alpha_{\text{tar}}\in\mathbb{C}$ (the encoding amplitude) of a single-mode coherent state $|\alpha_{\text{tar}}\rangle$ (the encoded state) from Alice to Bob following the operational principles in Sec. II.2. More details on coherent states can be found in Sec. S-II of [17]. When large coherent states are used in classical communications, we assume that Alice uses the simplest Binary Phase-shift keying (BPSK) modulation by mapping each of her classical bits according to a classical alphabet $\alpha_{\text{tar}}\in\{-\alpha,+\alpha\}$ with $\alpha\in\mathbb{R}$ being a large number. Specifically, $\alpha_{\text{tar}}$ is drawn from a distribution whose PDF is given by
@@ IV.1 Quantum State Evolution
We use the Characteristic Function (CF) formalism to describe the quantum state evolution in the process of coherent-state transfer (more details can be found in [17]). In coherent-state transfer without using diversity, the coherent state $\smash{|\alpha_{\text{tx}}\rangle=|\alpha_{\text{tar}}\rangle}$ is sent through a lossy atmospheric channel with transmissivity $T$ and excess noise $\epsilon$. Using the beam-splitter description of the atmospheric channel (see Appendix 2), we can express the CF of the output state $\rho^{(1)}$ that reaches the satellite receiver as
@@
When $\smash{M\geq 2}$ subchannels are used in diversity-assisted coherent transfer, the transmitted state in each subchannel becomes $|\alpha_{\text{tx}}^{(M)}\rangle=|\alpha_{\text{tar}}/\sqrt{M}\rangle$ due to Alice’s equal splitting (recall Sec. II.2). After the transmission over subchannels and Bob’s diversity combining based on the configuration described in Appendix C, the CF of Bob’s final state can be readily expressed using the beam-splitter description detailed in Appendix B.
For illustration purposes, in the following we explicitly present the CF of Bob’s final combined state in $\smash{M\times M}$ diversity-assisted coherent-state transfer with $M$ ranging from 2 to 4. When $M=2$, the combining module consists of only one beam splitter (see Fig. A3(a) in Appendix C), and the CF of the final state $\rho_{\text{DC}}^{(2)}$ is given by
@@
When $M=3$, the combining module consists of two beam splitters (see Fig. A3(b) in Appendix C), and the CF of the final state $\rho_{\text{DC}}^{(3)}$ is given by
@@
When $M=4$, the combining module consists of three beam splitters (see Fig. A3(c) in Appendix C), and the CF of the final state $\rho_{\text{DC}}^{(4)}$ is given by
@@ IV.2 Performance Metric
where $\chi_{1}(\xi)$ ($\chi_{2}(\xi)$) denotes the CF of $\rho_{1}$ ($\rho_{2}$[51]. Naturally, the effectiveness of transferring any given encoded state $|\alpha_{\text{tar}}\rangle$ from Alice to Bob can be quantified by the fidelity between Bob’s final state and a scaled version of Alice’s encoded state $\smash{|\alpha_{\text{tar}}^{\prime}\rangle=\left|{\alpha_{\text{tar}}/{\langle\sqrt{T}\rangle}}\right\rangle}$. Specifically, the scaling factor $\smash{{1}/{\langle\sqrt{T}\rangle}}$ is applied to reflect the necessary operation in a real-world system – Bob has to scale his measurement results (using the available information of channel statistics) to compensate for the (fixed) mean channel loss. Under one realization of the $M$ fluctuating subchannel transmissivities, such a fidelity can be calculated as $\smash{\mathcal{F}\left(\rho_{\text{DC}}^{(M)},|\alpha_{\text{tar}}^{\prime}\rangle\langle\alpha_{\text{tar}}^{\prime}|\right)}$ using Eq. (21). The analytical expression of this fidelity is presented in Appendix D.
@@ V.1 Entanglement Distribution
In Fig. 2 we adopt all the settings in Appendix 3 and plot the scaled logarithmic negativity $E_{\text{LN}}^{(M)}$ achieved by diversity-assisted entanglement distribution under different zenith angles (${\theta_{\text{z}}=0^{\circ}}$, ${\theta_{\text{z}}=30^{\circ}}$, and ${\theta_{\text{z}}=45^{\circ}}$) in a real-world uplink quantum communication system. From Fig. 2, we first observe that the TMSV state with a larger quadrature variance $V_{\text{s}}$ (which equivalently indicates a stronger squeezing level) is more susceptible to fading – such an observation is consistent with the findings in previous works, (e.g., [52]). It can also be clearly seen that the incorporation of diversity improves the effectiveness of entanglement distribution, and the use of more subchannels leads to performance improvement in terms of logarithm negativity. However, the performance improvement provided by adding subchannels is reduced as the number of subchannels increases. This “diminishing return” can be explained by Eq. (8) – indeed, when $M$ increases above a certain threshold, the non-Gaussian noise term $\frac{1}{M}\operatorname{Var}[\sqrt{T}](V_{\text{s}}-1)$ becomes negligible, making the total noise dominated by the term $(\langle\epsilon\rangle+1-T_{\text{eff}})$, which cannot be reduced by the use of diversity. Furthermore, the results in Fig. 2 also indicate that using diversity enables entanglement distribution over subchannels that are entanglement-breaking by themselves. For example, from Fig. 2 we can see that when $V_{\text{s}}=9$, it is impossible to distribute entanglement without using diversity under zenith angle $\smash{\theta_{\text{z}}=30^{\circ}}$ (see the $\smash{M=1}$ data point of the black dashed curve); however, the use of $2\times 2$ diversity enables the distribution of non-zero entanglement under the same channel condition (see the $\smash{M=2}$ data point of the black dashed curve).
[…]
The RCI not only gives a lower bound for the distillable entanglement but also represents a lower bound for the achievable secret key rate when the distributed entanglement is used for QKD. In CV-QKD, the modulation variance $V_{\text{mod}}$ (in a PM implementation), or, equivalently, the corresponding TMSV-state quadrature variance $V_{\text{s}}$ (in an EB implementation), is typically considered a free parameter chosen by Alice and Bob. Although the RCI is an information-theoretic metric that does not necessarily indicate the performance of a specific quantum information protocol (recall that the distributed entanglement can be used for purposes other than QKD), given the relevance of the RCI to QKD, one would intuitively be interested in optimizing $V_{\text{s}}$ for a maximized RCI. In Fig. 3, we adopt all the settings in Appendix 3 and plot the RCI $\mathcal{R}_{\text{C}}^{(M)}$ achieved by diversity-assisted entanglement distribution under different zenith angles (${\theta_{\text{z}}=0^{\circ}}$, ${\theta_{\text{z}}=30^{\circ}}$, and ${\theta_{\text{z}}=45^{\circ}}$) in a real-world uplink quantum communication system. Although the main focus of this work is not the performance optimization for any specific quantum information protocol (that utilizes the distributed entanglement), instead of plotting the RCI against the discrete number of subchannels $M$ (as we did for the logarithmic negativity in Fig. 2), in Fig. 3 we plot the RCI against the TMSV-state quadrature variance $V_{\text{s}}$, providing more useful insights by revealing the relationship between $V_{\text{s}}$ and the RCI in diversity-assisted entanglement distribution.
[…]
From Fig. 3, we first observe that most of the findings from Fig. 2 (related to the logarithmic negativity) also apply to the RCI, although some may be less apparent since the RCI is plotted against a different variable. Most importantly, it can be clearly seen that the use of diversity can improve the effectiveness of entanglement distribution in terms of the RCI – indeed, for a given $V_{\text{s}}$, the use of more subchannels (i.e., a larger $M$) can lead to a higher RCI under a given zenith angle $\theta_{\text{z}}$. On the contrary, without using diversity (i.e., when $M=1$), positive RCI cannot be achieved under any considered channel condition, regardless of the quadrature variance of the initial TMSV state $V_{\text{s}}$ (not shown).
Fig. 3 also reveals that under a given zenith angle $\theta_{\text{z}}$, there exists an optimal quadrature variance $V_{\text{s}}$ that gives the maximum RCI within the viable range of $V_{\text{s}}$ (in which a positive RCI can be achieved) for a given number of subchannels $M$. This can be understood with reference to the definition of the RCI in Eq. (13). Specifically, increasing the quadrature variance $V_{\text{s}}$ increases both $H(\rho_{\text{A}})$ and $H(\rho_{\text{AB}})$, but in different ways, resulting in the concave shape observed in the plotted curves. In general, the viable range of $V_{\text{s}}$ becomes narrower, and the maximum RCI becomes smaller under a larger $\theta_{\text{z}}$ due to the higher mean loss and stronger fading encountered, but the use of more subchannels broadens the viable range of $V_{\text{s}}$ and allows for a higher maximum RCI to be achieved. For example, under ${\theta_{\text{z}}=30^{\circ}}$, no positive RCI can be achieved when ${M=4}$ subchannels are used (indicated by the absence of the red dashed curve); however, positive RCI can be achieved when ${M=6}$ subchannels are used under the same channel condition (indicated by the green dashed curve). Similar to the diversity-enabled improvement in the logarithmic negativity, the diversity-enabled improvement in the RCI is also attributed to the $M$-time reduction of the fluctuation-induced non-Gaussian noise $\frac{1}{M}\operatorname{Var}[\sqrt{T}]\left(V_{\mathrm{s}}-1\right)$.
Here, we highlight the difference between our $\smash{M\times M}$ diversity-assisted entanglement distribution and the multiplexing of entanglement distribution over $M$ subchannels. When used alone, an entanglement-breaking channel cannot support effective entanglement distribution since Alice and Bob will end up sharing a separable state that does not contain any entanglement. Suppose the entanglement distribution is simply multiplexed over $M$ independent subchannels that are statistically identical to an entanglement-breaking channel. In that case, the total amount of distributed entanglement will still be zero since Alice and Bob will end up sharing $M$ separable states. As a result of the $M$-time reduction of the fluctuation-induced non-Gaussian noise, the unique performance enhancement that allows for entanglement distribution over entanglement-breaking subchannels is enabled by the use of $M\times M$ diversity and cannot be achieved by any spatial multiplexing scheme known to date.
@@ V.2 Coherent-State Transfer
We now examine the effectiveness of incorporating diversity for the purpose of coherent-state transfer within the context of SQCC. In addition to fading, the high mean loss over Earth-to-satellite channels can be very challenging for the faithful transfer of coherent states – indeed, our simulations indicate that the mean loss over a real-world Earth-to-satellite channel can be as high as ${\sim 40\,\text{dB}}$ (see Fig. A1 and Table A1 in Appendix 2 for details). In our investigation, we find that such a high mean loss essentially renders the faithful transfer of coherent states non-viable. It should be noted that this observation does not necessarily negate the possibility of classical communications, which focus on discriminating one coherent state from the other rather than the exact transfer of coherent states. Nevertheless, as in the case of coherent-state transfer, such a high loss will likely restrict the viability of the diversity-assisted discrimination of coherent states over fading channels (which has been investigated in [35]) as well. In order to demonstrate the usefulness of diversity as a means of fading mitigation in coherent-state transfer over other generic fading channels, we restrict ourselves to a regime with low mean subchannel losses and adopt a simple and flexible log-normal distribution (detailed in Appendix 4) to investigate the general behaviours of our diversity-assisted system over fading channels. Our focus on such a low-mean-loss regime makes our results (on average fidelity) more likely to be relevant to other FSO fading channels, such as those considered in [35] or satellite-based downlink channels that have a low mean loss and insignificant fading.
@@
[…]
@@
[…]
@@ VI Discussions
We have confirmed the usefulness of diversity for both entanglement distribution and coherent-state transfer over fading channels. Most straightforwardly, our findings indicate that employing diversity as a fading mitigation method can enhance the performance of all Gaussian quantum information protocols (including the standard CV-QKD) over fading channels (including Earth-to-satellite channels). As mentioned in Sec. II.1, in this work we have assumed that only the statistics of the transmissivity is known – in other words, the instantaneous channel transmissivity, sometimes referred to as the full Channel State Information (CSI), is not available to Bob. Interestingly, when the full CSI is available at Bob, we find that the diversity-assist system described in Sec. II.2 does not provide any non-negligible improvement in both entanglement distribution and coherent-state transfer under all our considered scenarios (results not presented).
@@
For a fair comparison across scenarios where different numbers of subchannels are used, we have assumed a fixed size for all receiving apertures, ensuring that the total received signal power at the satellite remains statistically the same regardless of the number of subchannels used. In addition to employing diversity (the effectiveness of which has been confirmed in this work), other methods for fading mitigation may also be beneficial in Earth-to-satellite CV quantum communications. The most intuitive alternative is the use of a single larger receiving aperture (despite the practical infeasibility of aperture averaging at the satellite). Considering the limited space available for a satellite, future research could involve comparing a single large receiving aperture with multiple smaller receiving apertures (occupying the same total area as the large receiving aperture) on the satellite and investigating how many smaller apertures (and of what size) are required to outperform a single large aperture.
It is worth noting that although the large coherence length at the satellite effectively negates the practical usefulness of engineering receiver-aperture separation at a single satellite, it is possible to exploit an additional degree of diversity and/or to achieve aperture averaging via the use of a distributed satellite constellation (where satellites are separated by large distances). Within this context, another useful future study may involve a detailed investigation into the formation and coordination of a satellite constellation for exploiting such diversity via appropriate inter-satellite distances.
@@ VII Conclusions
In this work, we explored the feasibility and effectiveness of utilizing spatial diversity as a means of fading mitigation in Earth-to-satellite CV quantum communications. We first determined a practical system model where spatial diversity can potentially be introduced to improve system performance within the context of satellite-based CV quantum communications. Based on our system model, we proposed a practical diversity-assisted scheme for the tasks of entanglement distribution and coherent-state transfer, and we analyzed the evolution of quantum states within these tasks. Using our numerical simulations for the realistic modelling of Earth-to-satellite channels, we perform a detailed performance investigation on diversity-assisted entanglement distribution, confirming the usefulness of diversity in Earth-to-satellite quantum communications. Within the context of SQCC over fading channels, our further results indicate that the transfer of both large (classically-modulated) coherent states (for the classical part of communications) and small (quantum-modulated) coherent states (for the quantum part of communications) can benefit from the use of diversity. The usefulness of the uplink configuration in space-based quantum communications is further elaborated on in Sec. S-III of [17]. Note that our results can be applied to arbitrary FSO fading channels.